OpenStudy (anonymous):

The distance between the points (1,8) and (4,4) is equal to 5

OpenStudy (anonymous):

TRUE OR FALSE

OpenStudy (accessdenied):

\[ D = \sqrt{(x_2 - x_1)^2 - (y_2 - y_1)^2} sqrt((4-1)^2 + (4-8)^2) = sqrt((3)^2 + (-4)^2) = sqrt(9 + 16) = sqrt(25) = 5 true

OpenStudy (accessdenied):

\[ D = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2} \] maybe this would look better with a plus sign too

OpenStudy (anonymous):

TY

OpenStudy (accessdenied):

Glad to help. :)